solve: 2/x+1 + 3/x-1 = 5/x^2-1
Answer by stanbon(75887)

solve: 2/x+1 + 3/x-1 = 5/x^2-1
---
Multiply thru by x^2-1 to get:
2(x-1) + 3(x+1) = 5
5x+1 = 5
5x = 4
x = 4/5
============
Cheers,
Stan H.